OpenStudy (anonymous):

find the exact value of: sin13pi/8

OpenStudy (anonymous):

A) -sqrt2+sqrt2/4 B)sqrt2+sqrt2/4 C)-sqrt2-sqrt2/4 D)sqrt2-sqrt2/4

OpenStudy (anonymous):

13pi/8=2pi-3pi/8

OpenStudy (anonymous):

sin(13pi/8)=-sin(3pi/8)=-sqrt((1-cos(3pi/4))/2)

OpenStudy (anonymous):

which one do you think it is? a, b, c, or d?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

\[-\sqrt{(2+\sqrt{2})/4}\]

OpenStudy (anonymous):

thank yoooooooouuuuuuuuu :)